OpenStudy (anonymous):

solve for x 3(x-1)=2x

OpenStudy (anonymous):

3x-3=2x 3x-2x=3 x=3

OpenStudy (anonymous):

FIRST, DISTRIBUTE 3 to (x-1) you will get 3x-3=2x, then transpose 2x and 3 3x-2x=3, combine like terms, x=3 so, the value of x is 3. check it by substituting 3 to x.